OpenStudy (anonymous):

Starting from rest , an airplane (mass= 12,000 kg) accelerates 1,700 meters along a runway . What is the average net force on the plane if its final speed is 55 m/s ?

OpenStudy (anonymous):

F = m*a

OpenStudy (anonymous):

speed v depends on acceleration over time there is no time given

OpenStudy (jt950):

v^2=u^2 + 2 a s u=0 so, 55^2=2a(1700) a=0.94 F=ma=(12000)(0.94)=11,318.91N (Correct to 2 decimal place)